OpenStudy (anonymous):

Please help I don't think there is a way to solve this problem. I don't think there is enough info. Thanks for your time A box with a square base has a volume of 12 cubic feet. How high should the box be in order that it has the minimal surface area. has a top and 6 sides

OpenStudy (anonymous):

please please help

OpenStudy (amistre64):

so it looks like we need to define a volume of a box that is 12 cubic feet ... and work some derivatives

OpenStudy (amistre64):

do you know a general equation for surface area of a box with a square base?

OpenStudy (amistre64):



OpenStudy (amistre64):

Sa = 4sh + 2s^2 V = s^2 h = 12 h = 12/s^2 Sa = 4s(12/s^2) + 2s^2 = 48s^(-1) + 2s^2 when we know s, we should then be able to solve for h

OpenStudy (anonymous):

Is the right answer for h 0.0833333

OpenStudy (amistre64):

dunno, i havent had the chance to work it out :) ill check on it when i can leter.